# Revisiting Lesion Tracking in 3D Total Body Photography

### v1

@@ 1 Introduction
Total body photography (TBP) captures the entire body of a patient using 2D images [19] and/or a 3D mesh [44, 16]. As such, TBP can be an effective for monitoring the evolution of lesions [9, 20, 23, 18, 43, 42]. From a systematic review, a large number (>100) of common nevi is required for investigation by TBP in most studies [20].
@@ 2.1 Shape correspondence for humans
Shape correspondence between non-rigid surfaces represented as triangle meshes has been an active research topic in computer vision and computer graphics [57, 46, 10]. Shape correspondence provides dense correspondence (e.g. vertex-to-vertex or point-to-point) across 3D meshes. For human shapes, priors of human body are commonly applied, such as near-isometric deformation and local rigidity for limbs [65, 61, 5, 4].
@@ 2.2 Graph matching
Graph matching is a fundamental problem in computer vision and pattern recognition that involves finding a correspondence between the nodes and edges of two graphs. Two-graph matching can be modeled as the quadratic assignment problem (QAP), and is known to be NP-hard.
@@
Learning-based methods have been shown to improve matching accuracy [60, 45, 30]. Wang et al. [60] presents a QAP network directly in which the matching problem is translated into a constrained vertex classification task. Liao et al. [30] convert the problem of hypergraph matching into a node classification problem and develop a hypergraph neural network. Despite their promise, these methods often require exhaustively annotated datasets for training and struggle to generalize across different domains or datasets. Meanwhile, to address real-world scenarios involving noisy or incomplete data, some techniques are developed for partial graph matching and soft assignment [58, 15].
@@ 2.3 Skin lesion tracking in total body photography
Recently, the concept of finding lesion correspondence using a 3D representation of the human body has been explored in [7, 64, 3, 21]. Zhao et al. [64], Bogo et al. [7], and Ahmedt-Aristizabal et al. [3] proposed to use a template mesh and rely on anatomical position defined on the template mesh for lesion matching. However, using the template-based correspondence map is insufficient to accurately match lesions. Furthermore, these methods have limited positional expression for lesions since they “snap” the location of a lesion to the nearest vertex. Huang et al. [21] proposed to improve the lesion correspondence localization accuracy using landmark-based correspondences refined by texture information. However, their method requires manual annotation of landmarks, is limited to the resolution of the mesh, and is sensitive to inconsistent texture from scan to scan. In this paper, we represent lesion positions using barycentric coordinates within a triangle, making our approach less sensitive to the resolution of the mesh and allowing us to achieve a higher matching accuracy. Additionally, we utilize lesion signals that are agnostic to the textured mesh.
@@ 3.1 Problem Formulation
Given a template mesh $\mathcal{M}_{\mathcal{T}}$, source and target meshes $\mathcal{M}_{0}$, $\mathcal{M}_{1}$, and two sets of detected lesions $X_{0}\subset\mathcal{M}_{0}$, $X_{1}\subset\mathcal{M}_{1}$, we would like to find correspondence maps $\phi_{0}^{\mathcal{T}}:\mathcal{M}_{0}\rightarrow\mathcal{M}_{\mathcal{T}}$ and $\phi_{1}^{\mathcal{T}}:\mathcal{M}_{1}\rightarrow\mathcal{M}_{\mathcal{T}}$ such that identical lesions are mapped to the same position in the template mesh $\mathcal{M}_{\mathcal{T}}$. We also solve for a matching matrix $\pi=\{0,1\}^{(|X_{0}|+1)\times(|X_{1}|+1)}$. By adding a dummy lesion to each of the lesion sets, we allow the matching function to account for unmatchable lesions. Specifically, assuming a lesion in the source scan can be matched to at most one lesion in the target scan and vice versa, we enforce:
@@ 3.2.1 Template-based coarse correspondence
We start by constructing a coarse correspondence map between the source/target and a template mesh. We follow the approach from Marin et al. [34], denoted by SMPL-NICP hereafter. Let $\mathcal{M}_{\mathcal{T}}$ be the template mesh, for an input mesh $\mathcal{M}_{i},\ i=\{0,1\}$, the output from SMPL-NICP is a deformed mesh $\mathcal{M}_{\mathcal{T}_{i}}$ registered to $\mathcal{M}_{i}$. To construct a correspondence map $\phi_{i}^{\mathcal{T}}:\mathcal{M}_{i}\rightarrow\mathcal{M}_{\mathcal{T}}$, we first register the template mesh to $\mathcal{M}_{i}$ with $\mathcal{M}_{\mathcal{T}_{i}}$. We note that $\mathcal{M}_{\mathcal{T}_{i}}$ and $\mathcal{M}_{\mathcal{T}}$ share the same topology. We define $\psi_{\mathcal{T}_{i}}^{\mathcal{T}}:\mathcal{M}_{\mathcal{T}_{i}}\rightarrow{\mathcal{M}_{\mathcal{T}}}$ as an operation relying on topology to take a point in the space of the deformed mesh to the space of the template mesh in the canonical pose. Therefore, for a point $p_{i}\in\mathcal{M}_{i}$, a correspondence map $\phi_{i}^{\mathcal{T}}:\mathcal{M}_{i}\rightarrow\mathcal{M}_{\mathcal{T}}$ can be defined as:
$\phi_{i}^{\mathcal{T}}(p_{i})=\psi_{\mathcal{T}_{i}}^{\mathcal{T}}\Bigl{(}\operatorname*{arg\,min}_{q\in\mathcal{M}_{\mathcal{T}_{i}}}||q-p_{i}||\Bigr{)}$ (3)
Reversely, we can construct a correspondence map $\phi_{\mathcal{T}}^{i}:\mathcal{M}_{\mathcal{T}}\rightarrow\mathcal{M}_{i}$ by:
$\phi_{\mathcal{T}}^{i}(q)=\operatorname*{arg\,min}_{p_{i}\in\mathcal{M}_{i}}||p_{i}-\psi_{\mathcal{T}}^{\mathcal{T}^{i}}(q)||,$ (4)
where $\psi_{\mathcal{T}}^{\mathcal{T}^{i}}:\mathcal{M}_{\mathcal{T}}\rightarrow\mathcal{M}_{\mathcal{T}_{i}}$ brings a point from the canonical space to the deformed mesh. Fig. 1 illustrates the template mesh in (a), the source and the target meshes in (b), the deformed template meshes registered to the source and the target in (c), the source and the target lesions found in the deformed meshes in (d), and the source and the target lesions mapped to the template mesh in (e).
@@ 3.2.2 Surface point correspondence map
Given a vertex-to-surface-point correspondence map $\Phi_{i}^{j}:\mathcal{V}_{i}\rightarrow\mathcal{M}_{j}$, we use the barycentric encoding to extend it to a surface-point-to-surface-point correspondence map $\phi_{i}^{j}:\mathcal{M}_{i}\rightarrow\mathcal{M}_{j}$. Concretely, to find the correspondence of a surface point $p\in\mathcal{M}_{i}$ to the mesh $\mathcal{M}_{j}$, we map the three vertices of the triangle containing the point $p$ onto the mesh $\mathcal{M}_{j}$, interpolate the positions of the imaged vertices using the barycentric coordinate of $p$, and then find the point on $\mathcal{M}_{j}$ closest to the interpolant. This operation is accurate in practice if the three imaged vertices are close. Formally, for a point $p\leftrightarrow(t^{p},\{\alpha^{p},\beta^{p},\gamma^{p}\})\in\mathcal{M}_{i}$ with $\mathcal{F}_{i}(t^{p})=(v_{0}^{p},v_{1}^{p},v_{2}^{p})$ representing the triangle containing $p$, we have:
@@
Figure 1: Visualization of the source and target lesions mapped to a template mesh by registering the template mesh to the source and target meshes. (a) shows the template mesh $\mathcal{M}_{\mathcal{T}}$. (b) shows the source mesh ($\mathcal{M}_{0}$) and the target mesh ($\mathcal{M}_{1}$). (c) shows the deformed template meshes, $\mathcal{M}_{\mathcal{T}_{0}}$ and $\mathcal{M}_{\mathcal{T}_{1}}$ that register to the source and the target, respectively. (d) visualizes the deformed meshes along with the lesions mapped to the deformed meshes. (e) shows the source and the target lesions mapped to the template mesh. Lesions in correspondences are in the same color.
@@ 3.3.2 Surface optical flow
with the first and the second terms penalizing the failure of the vector field to explain the difference in the texture signal and the lesion signal respectively, the third term encouraging the smoothness of the flow, and the fourth term regularizing the norm of the flow to respect the initial correspondence map. We follow the approach proposed by Prada et al., solving for the flow field $\vec{v}$ hierarchically. Please refer to [41] for more details. Fig. 2 visualizes the source and the target (a) texture and (b) lesion signals transferred to the template mesh. Fig. 2 (c) shows the vector field obtained by minimizing Eqn.6.
@@ 3.4 Lesion Assignment
where $D_{\mathcal{T}}:\mathcal{M}_{\mathcal{T}}\times\mathcal{M}_{\mathcal{T}}\rightarrow{\mathbb{R}}^{\geq 0}$ is the geodesic distance function on $\mathcal{M}_{\mathcal{T}}$. The assignment problem can be optimized through the Kuhn–Munkres algorithms [38]. We follow the implementation in Pygmtools [59] to solve the minimizer to Eqn. 9.

### v2

@@ 1 Introduction
Total body photography (TBP) captures the entire body of a patient using 2D images [20] and/or a 3D mesh [45, 17]. As such, TBP can be effective for monitoring the evolution of lesions [10, 21, 24, 19, 44, 43]. From a systematic review, an individual with a large number (>100) of common nevi is usually included in the target population of TBP [21].
@@ 2.1 Shape correspondence for humans
Shape correspondence between non-rigid surfaces represented as triangle meshes has been an active research topic in computer vision and computer graphics [58, 47, 11]. The shape correspondence problem for triangle meshes is finding a set of corresponding points between two meshes. For human shapes, priors of the human body are commonly applied, such as near-isometric deformation and local rigidity for limbs [66, 62, 6, 4].
@@ 2.2 Graph matching
Given a set of lesions detected in a mesh, we can construct a graph in which a node corresponds to a single lesion and an edge is a connection between a pair of lesions. The node attribute is the position of a lesion, and the edge attribute is the geodesic distance between a pair of lesions. Then, the problem of matching source and target lesions can be formulated as a (partial) graph-matching problem that maximizes the node-to-node and edge-to-edge affinity of the two graphs. Two-graph matching can be modeled as the quadratic assignment problem (QAP), and is known to be NP-hard.
@@
Learning-based methods have been shown to improve matching accuracy [61, 46, 31]. Wang et al. [61] presents a QAP network to solve the matching problem as a vertex classification task over the association graph whose nodes represent candidate correspondences between the two graphs and edge weights are induced by the affinity matrix built with the two graphs. Liao et al. [31] convert the problem of hypergraph matching into a node classification problem and develop a hypergraph neural network. Despite their promise, these methods often require exhaustively annotated datasets for training and struggle to generalize across different domains or datasets. Meanwhile, to address real-world scenarios involving noisy or incomplete data, some techniques are developed for partial graph matching and soft assignment [59, 16].
@@ 2.3 Skin lesion tracking in total body photography
Recently, the concept of finding lesion correspondence using a 3D representation of the human body has been explored in [8, 65, 3, 22]. Zhao et al. [65], Bogo et al. [8], and Ahmedt-Aristizabal et al. [3] proposed to use a template mesh and rely on anatomical position defined on the template mesh for lesion matching. However, using the template-based correspondence map is insufficient to accurately match lesions. Furthermore, these methods have limited positional expression for lesions since they “snap” the location of a lesion to the nearest vertex. Huang et al. [22] proposed to improve the lesion correspondence localization accuracy using landmark-based correspondences refined by texture information. However, their method requires manual annotation of landmarks, is limited to the resolution of the mesh, and is sensitive to inconsistent texture between scans due to scanning artifacts. In this paper, we represent lesion positions using barycentric coordinates within a triangle, making our approach less sensitive to the resolution of the mesh and allowing us to achieve a higher matching accuracy. Additionally, we utilize lesion signals that are agnostic to the textured mesh assuming lesions are provided separately.
@@ 3.1 Problem Formulation
Given a template mesh $\mathcal{M}_{\mathcal{T}}$, source and target meshes $\mathcal{M}_{0}$, $\mathcal{M}_{1}$, and two sets of detected lesions $X_{0}\subset\mathcal{M}_{0}$, $X_{1}\subset\mathcal{M}_{1}$, we would like to find correspondence maps $\phi_{0}^{\mathcal{T}}:\mathcal{M}_{0}\rightarrow\mathcal{M}_{\mathcal{T}}$ and $\phi_{1}^{\mathcal{T}}:\mathcal{M}_{1}\rightarrow\mathcal{M}_{\mathcal{T}}$, and a matching matrix $\pi=\{0,1\}^{(|X_{0}|+1)\times(|X_{1}|+1)}$ minimizing an energy of consisting of two terms:
$E_{X_{0},X_{1}}(\phi_{0}^{\mathcal{T}},\phi_{1}^{\mathcal{T}},\pi)=\sum_{X_{0},X_{1}}E_{\text{DistanceProximity}}(\phi_{0}^{\mathcal{T}},\phi_{1}^{\mathcal{T}},\pi)+E_{\text{Stochasticity}}(\pi)\ .$ (1)
That is, we want a pair of corresponding source and target lesions to be close to each other while encouraging the correspondence matrix to be doubly stochastic. By adding a dummy lesion to each of the lesion sets in $\pi$, we allow the matching function to account for unmatchable lesions. Specifically, assuming a lesion in the source scan can be matched to at most one lesion in the target scan and vice versa, we also enforce:
@@ 3.2.1 Template-based coarse correspondence
We start by constructing a coarse correspondence map between the source/target and a template mesh. We follow the approach from Marin et al. [35] to acquire a deformed template mesh registered to the source/target mesh that allows us to construct the correspondence map. Given an input mesh, they propose a localized neural fields network in which a neural field is dedicated to a local region of body shape to predict the vertex displacement of the template mesh (SMPL [33] model). The parameters of the neural field are then refined using Iterative Closest Point [5] through backpropagation. Then, the updated neural field is utilized to register the SMPL model to the input, followed by a refinement that optimizes Chamfer distance. We denote the method SMPL-NICP. Let $\mathcal{M}_{\mathcal{T}}$ be the template mesh, for an input mesh $\mathcal{M}_{i},\ i=\{0,1\}$, the output from SMPL-NICP is a deformed template mesh (i.e. with the same topology as the original template) whose geometry is registered to that of $\mathcal{M}_{i}$.
We define a correspondence map $\phi_{i}^{\mathcal{T}}:\mathcal{M}_{i}\rightarrow\mathcal{M}_{\mathcal{T}}$, by first deforming the template mesh to $\mathcal{M}_{i}$ and then finding, for every point $p\in\mathcal{M}_{i}$, the nearest surface point on the deformed template. Similarly, we construct a correspondence map $\phi_{\mathcal{T}}^{i}:\mathcal{M}_{\mathcal{T}}\rightarrow\mathcal{M}_{i}$ by finding the closest surface point on the input mesh $\mathcal{M}_{i}$ for each point on the deformed template mesh. We note that $\phi_{i}^{\mathcal{T}}$ and $\phi_{\mathcal{T}}^{i}$ are not inverses of each other since two different points on the source/target can have the same closest point on the deformed template. Fig. 1 illustrates the template mesh in (a), the source and the target meshes in (b), the correspondence maps from the source/target to the template in (c), and the source and the target lesions mapped to the template mesh in (d).
@@ 3.2.2 Surface point correspondence map
Given a vertex-to-surface-point correspondence map $\Phi_{i}^{j}:\mathcal{V}_{i}\rightarrow\mathcal{M}_{j}$, we use the barycentric encoding to extend it to a surface-point-to-surface-point correspondence map $\phi_{i}^{j}:\mathcal{M}_{i}\rightarrow\mathcal{M}_{j}$. Concretely, to find the correspondence of a surface point $p\in\mathcal{M}_{i}$ to the mesh $\mathcal{M}_{j}$, we map the three vertices of the triangle containing the point $p$ onto the mesh $\mathcal{M}_{j}$, interpolate the positions of the imaged vertices using the barycentric coordinate of $p$, and then find the point on $\mathcal{M}_{j}$ closest to the interpolant. Formally, for a point $p\leftrightarrow(t^{p},\{\alpha^{p},\beta^{p},\gamma^{p}\})\in\mathcal{M}_{i}$ with $\mathcal{F}_{i}(t^{p})=(v_{0}^{p},v_{1}^{p},v_{2}^{p})$ representing the triangle containing $p$, we have:
@@
Figure 1: Visualization of the source and target lesions mapped to a template mesh by registering the template mesh to the source and target meshes. (a) shows the template mesh $\mathcal{M}_{\mathcal{T}}$. (b) shows the source mesh ($\mathcal{M}_{0}$) and the target mesh ($\mathcal{M}_{1}$). (c) shows the correspondence maps from the source/target to the template. (d) shows the source and the target lesions mapped to the template mesh. Lesions in correspondences are visualized in the same color.
@@ 3.3.2 Surface optical flow
with the first and the second terms penalizing the failure of the vector field to explain the difference in the texture signal and the lesion signal respectively, the third term encouraging the smoothness of the flow, and the fourth term regularizing the norm of the flow to respect the initial correspondence map. We follow the approach proposed by Prada et al., solving for the flow field $\vec{v}$ hierarchically. Please refer to [42] for more details. Fig. 2 visualizes the source and the target (a) texture and (b) lesion signals transferred to the template mesh. Fig. 2 (c) shows the vector field obtained by minimizing Eqn.5.
@@ 3.4 Lesion Assignment
where $D_{\mathcal{T}}:\mathcal{M}_{\mathcal{T}}\times\mathcal{M}_{\mathcal{T}}\rightarrow{\mathbb{R}}^{\geq 0}$ is the geodesic distance function on $\mathcal{M}_{\mathcal{T}}$. The assignment problem can be optimized through the Kuhn–Munkres algorithms [39]. We follow the implementation in Pygmtools [60] to solve the minimizer to Eqn. 8.